# Investigating the Origin of CMB Large-Scale Features Using LiteBIRD and CMB-S4

### v1

@@ 1 Introduction
Primordial features are examples of departures in the potential or kinetic term from SR that induce strongly scale-dependent components in the primordial density perturbations. Besides revealing the shape of the inflaton potential (e.g. [19, 20, 21, 22]), primordial features can momentarily increase the energy scale of inflation, potentially boosting the energy reach [23] of this “cosmological collider” [24, 25] by orders of magnitude, allowing for exploration of fundamental physics at even higher energies beyond what can be observed in ground-based collider experiments. Primordial features induced by oscillating massive fields could also be used to probe the background evolutionary history of the primordial universe, providing a possible form of direct evidence for the inflation scenario [26, 27, 28, 29]. (See [30, 31, 32, 33] for reviews on primordial features.) While there is no evidence for such features in current CMB observations, there are some notable, although not statistically significant, deviations from the concordance model at several multipoles in the measured angular spectra. In particular, deviations at very large scales in the $TT$ power spectrum have been observed by both WMAP and Planck. These so-called “$-\ell$ anomalies” are manifest in the $TT$ spectrum as a suppression of power at $\ell<10$ and a dip at $\ell\simeq 20$, the latter being approximately $2\sigma$ away from the $\Lambda$CDM best fit. These anomalies have been explored as possibly being primordial in origin [34, 20, 35, 36, 37, 38, 39]. While these anomalies may be explained by statistical fluctuations, especially when noting the large uncertainties at $\ell<10$ due to cosmic variance, future experiments may determine whether these deviations are significant, and they are therefore worth exploring.
@@
A slew of next generation CMB missions following Planck are under development. Some notable examples include LiteBIRD [40], Simons Observatory [41], and CMB-S4 [42]. Since Planck has achieved nearly cosmic-variance limited measurements in the temperature anisotropies up to $\ell\sim 2500$, most of the improvement in precision from these missions will be in polarization. LiteBIRD is planned for launch by the late-2020s and will cover satellite observations with a focus on polarization at large scales ($\ell\leq 200$). Simons Observatory and CMB-S4 are both ground-based experiments. Simons Observatory began taking data in mid-2024, probing multipoles up to $\ell=4000$. CMB-S4 observations are planned for the late-2020s, probing multipoles up to $\ell=5000$. One of the primary goals of these experiments is to obtain high-precision $B$-mode measurements in order to detect primordial gravitational waves. Since the amplitude of the $B$-mode anisotropies is small compared to that of $E$-mode and temperature anisotropies, these missions require unprecedented precision of $E$-mode measurements, with the $EE$ and $TE$ power spectra projected to being constrained nearly to the level of cosmic variance. (Planck measurements have already achieved precision near the level of cosmic variance in the $TT$ power spectrum.) Since models designed to explain the $-\ell$ anomalies in the $TT$ spectrum also predict correlated features in the $EE$ and $TE$ spectra, future experiments will provide a more accurate test of these models.
@@
In this paper, we test the ability of future CMB experiments to constrain or detect cosmological models beyond $\Lambda$CDM that modify the CMB power spectra at large scales with a particular emphasis on primordial feature models. Current candidates to explain the $-\ell$ anomalies include such primordial feature models, which modify the shape of the scalar primordial power spectrum from inflation, as well as models that alter the late-time evolution of the universe, affecting the integrated Sachs-Wolfe (ISW) effect. As a representative sample, we consider three popular models with distinct large-scale effects in the CMB power spectrum: (1) a primordial feature introduced as a step in the inflationary potential that can produce either a bump or a dip at large scales, used to explain the $\ell\simeq 20$ dip in $TT$; (2) a second primordial feature model introducing a kink in the potential, which can lead to either large-scale suppression or amplification, used to explain the $\ell<10$ suppression in $TT$; and (3) a model of dark energy parameterized by the equation of state $w$ of the dark energy fluid and the speed of sound $c_{s}^{2}$ of its perturbations, modifying the late-time ISW effect and also producing either large-scale suppression or amplification. For completeness, we also consider a fourth model, characterized by a micron-sized fifth spacetime dimension that predicts only amplification of the scalar primordial power spectrum, which is already disfavored by current observations. We use the fully Bayesian pipeline introduced in [43] and later improved in [44]. We first compare the models against real Planck data, providing constraints on their parameters, identifying best-fit candidates, and assessing their observational evidence compared to the standard cosmological model. We then perform a Bayesian forecast analysis for the $TT$, $EE$, and $TE$ power spectra using the projected sensitivities to temperature and $E$ modes from future experiments. While the main driver of our analysis is large-scale measurements from LiteBIRD, most models aiming to fit the $-\ell$ anomalies produce modifications to the anisotropy spectra at smaller scales as well. Therefore, we complement our forecast with projected smaller-scale measurements. While such high-quality anisotropy measurements will soon be available from the Simons Observatory, our purpose is to set the most optimistic prospects by including small-scale data, so we provide projected forecasts from CMB-S4. As a final test of LiteBIRD’s capabilities, we ignore the short-scale effects of the primordial feature models and only focus on the suppression and dip anomalies at large scales, which we represent with non-physical, diagnostic templates that produce no small-scale modifications to the primordial power spectrum. We repeat our data and forecast analyses on these templates. We aim to answer three major questions with our analysis: (1) If the true model of the universe is featureless (i.e., consistent with the concordance model), can we hope to rule out large-scale features as statistical fluctuations? (2) Conversely, if the true model contains some kind of large-scale feature, can we detect such a feature? (This question has also been studied for the step model in [39].) (3) In the case of (2), can we further discriminate between different types of features with future observations?
@@ 2 The Models
We begin by presenting an overview of the models used in this study. We consider three types of large-scale deviations from the concordance model: a dip, suppression, and amplification in the $TT$ power spectrum. Our goal is to investigate if future experiments are able to detect such deviations and discriminate between them.
@@
This model is convenient as it allows us to easily relate the free parameters of the model to the scalar power spectrum amplitude $A_{s}$ and tilt $n_{s}$ as well as to the tensor-to-scalar ratio $r$, all defined at a given scale with the Hubble crossing condition $k_{0}=a_{0}H_{0}$
@@
From these equations, we see that $\epsilon_{0}$ controls the amplitude of the tensor-to-scalar ratio $r$. Since our analysis in this paper does not make use of $B$-mode data, we follow previous works [43, 46, 47, 44] and fix $\epsilon_{0}=10^{-7}$ in order not to produce observable $B$ modes.
Dip at Large Scales. We first consider a sharp primordial feature introduced as a bump or dip in the primordial power spectrum. Such a feature can be generated by a “step” in the inflaton potential. This step model was first introduced by [20] and is characterized by the potential
@@
where $A$ is the height of the step, $\phi_{0}$ is the feature location, and $\Delta$ is the feature width. The initial perturbation in the power spectrum at the feature location can be described as either a large dip if $A>0$ (a downward step) or a large bump if $A<0$ (an upward step). Introducing such a feature is motivated by fitting the $\ell\simeq 20$ dip in the $TT$ angular power spectrum (e.g., see [48, 37, 49]). This model introduces three model parameters: $A$, $\phi_{0}$, and $\Delta$. We relate these model parameters to the following three “effective parameters,” which are more intuitive and more closely related to observables: the maximum deviation in the primordial power spectrum from the featureless spectrum $\Delta\mathcal{P}/\mathcal{P}_{0}$, the number of $e$-folds defined from the beginning of inflation ($N_{\rm start}\equiv 0$) at which the feature occurs $N_{0}$, and the feature duration in terms of number of $e$-folds $\Delta N$. The relations between the model and effective parameters are derived in Appendix A.
Large-Scale Primordial Suppression. Next, we consider a primordial feature model that results in large-scale suppression by introducing a “kink” in the inflaton potential. This model was first proposed by [19] and is defined by the potential
@@
where $A$ is the initial slope of the potential, $\phi_{0}$ is the feature location, and $\Delta$ is the width of the kink. This feature is characterized by an sharp change in the shape of the potential, which is of the form $V\sim V_{0}(\phi)[1+A(\phi_{0}-\phi)]$ for $\phi\ll\phi_{0}$ and $V\sim V_{0}$ for $\phi\gg\phi_{0}$. This model can result in either suppression in the power spectrum if $A>0$ (a negative initial slope) or amplification if $A<0$ (a positive initial slope). This model is therefore motivated by fitting the $\ell<10$ suppression in the $TT$ angular power spectrum. The same model parameters introduced in the step model also describe the kink model, and these are similarly related to the effective parameters $\Delta\mathcal{P}/\mathcal{P}_{0}$, $N_{0}$, and $\Delta N$. We note that other models exist that predict a similar suppression of power that could fit the $-\ell$ suppression observed in the temperature data. For example, double inflation models where inflation consists of two stages driven by two kinetically coupled scalar fields produces a similar feature signal to that of the kink model [50]. Other single field models such as “punctuated inflation” [51, 52], fast roll initial conditions on the inflaton [53, 54, 55, 56], or inflationary models with non-zero spatial curvature [57, 58] also produce suppression, although more drastic than that considered here.
@@
Finally, while our analysis focuses on the CMB and the ability of this model to fit the $-\ell$ anomalies, it is important to note that models with a time-evolving equation of state also significantly impact the background evolution of the universe and the late-time clustering of matter. Consequently, they are strongly constrained by other datasets, such as supernovae and LSS data. However, we do not include these datasets in our analysis. Therefore, the constraints on this model presented here should be considered conservative and are expected to become more stringent with the inclusion of these additional datasets.
@@
where $\epsilon$ and $\delta$ are respectively the first and second SR parameters. We have defined the variable $x\equiv\pi kR_{0}$, where $R_{0}$ is the compact dimension radius at the beginning of inflation. Since the initial value of $R_{0}$ is not fixed in [71], we convert it into a parameter to be determined by our analysis. We relate $R_{0}$ to the effective parameter $N_{\rm DD}-N_{*}$ via
@@
Figure 1: Examples of large-scale effects in the primordial, $TT$, $TE$, and $EE$ power spectra [from left to right] for each of the models considered in this paper. The Planck 2018 data are represented by the gray points. For these examples, the step model parameters are $\ln(10^{10}A_{s})=3.14$, $n_{s}=0.966$, $N_{0}-N_{*}=-4.0$, $\Delta N=0.5$, and $\Delta\mathcal{P}/\mathcal{P}_{0}=-0.6$. The kink model parameters are the same as those for the step model except for $\Delta N=0.1$ and $\Delta\mathcal{P}/\mathcal{P}_{0}=-0.5$. The featureless and $w-c_{s}^{2}$ models were calculated using the same parameters for the kink model but with $\Delta\mathcal{P}/\mathcal{P}_{0}=0$. The additional $w-c_{s}^{2}$ model parameters are $w=-0.5$ and $c_{s}^{2}=0.01$, with $H_{0}=53.0~{}{\rm km}~{}{\rm s}^{-1}~{}{\rm Mpc}^{-1}$. The dark dimension model parameters are $\ln(10^{10}A_{s})=3.04$, $n_{s}=0.966$, and $N_{\rm DD}-N_{*}=-4.0$ with $A_{s}$ and $n_{s}$ calculated at $k_{*}=0.05~{}{\rm Mpc^{-1}}$.
@@ 3 Methodology
We first constrain the effective parameters using the latest CMB temperature and $E$-mode polarization data from the Planck mission [1]. For the multipole range $\ell=2-29$, we use the commander_dx12_v3_2_29 (lowT) likelihood for the temperature anisotropies and the simall_100x143_offlike5_EE_Aplanck_B (lowE) likelihood for the $E$-mode anisotropies. For small angular scales $\ell=30-2500$, we use the 2020 CamSpec v12.5111https://people.ast.cam.ac.uk/s̃tg20/camspec/ (EG20) likelihood [72]. Because of the wide range of possible effective parameters that could be chosen given the CMB data, especially considering the large uncertainties at low $\ell$, we expect multimodal posterior distributions in our samples. Therefore, we use nested sampling to explore the parameter space using CosmoChord, a PolyChord [^2] https://github.com/PolyChord add-on of CosmoMC [^3] https://cosmologist.info/cosmomc/ [73, 74, 75].
@@
Our choice of ranges for the priors of the cosmological and effective parameters are listed in Table 1. We purposefully choose a prior volume large enough to account for a wide range of possible effects, including the possibility of a featureless spectrum so that the baseline model is also nested into the extended ones. As mentioned, we set the energy scale of inflation to $\epsilon\equiv 10^{-7}$. For the featureless, step, kink, and $w-c_{s}^{2}$ models, we calculate the primordial power spectrum numerically using BINGO [76]. By convention, we set the end of inflation to $N_{\rm end}=68$ and the pivot scale to exit the horizon at $N_{*}=N_{\rm end}-50=18$ $e$-folds from the onset of inflation. We also define $A_{s}$ and $n_{s}$ at the scale $k_{0}$, which exits the horizon $N_{0}$ $e$-folds from the start of inflation, corresponding to the feature scale for the step and kink models. For the featureless and $w-c_{s}^{2}$ models, we set $\Delta\mathcal{P}/\mathcal{P}_{0}=0$ so that $N_{0}-N_{*}$ and $\Delta N$ are arbitrary, as no features show up in $\mathcal{P}(k)$. We fix them to $N_{0}-N_{*}=-4.0$ and $\Delta N=1.0$. For the dark dimension model, the primordial power spectrum is evaluated directly using Eq. (2.9). For this model, $A_{s}$ and $n_{s}$ are defined at the pivot scale $k_{*}\equiv 0.05~{}{\rm Mpc}^{-1}$, which exits the horizon at $N_{*}=N_{\rm end}-50$. The compact dimension scale $k_{\rm DD}$ exits the horizon at $N_{\rm DD}$ $e$-folds from the onset of inflation. [^4] Note that the dark dimension scale $k_{\rm DD}$ is different from what we call the feature scale $k_{0}$ for the other primordial models we consider. Specifically, $k_{\rm DD}$ does not correspond to the point where the transition from $n_{s}\sim 0$ to $n_{s}\sim 1$ takes place, which one could associate with the feature scale $k_{0}$. Instead, we note that $k_{0}\approx 5k_{\rm DD}$, or $N_{0}-N_{*}=\ln 5+N_{\rm DD}-N_{*}$, which arises from the $\coth x+x{\rm csch}^{2}x$ term in Eq. (2.9). We choose our prior range for $N_{\rm DD}-N_{*}$ such that $N_{0}-N_{*}$ in all of our primordial models has the same prior range.
To determine the improvement in the likelihood of one model over another, we report the Bayes factor for each model, defined as the ratio of evidences between two models. We calculate the Bayes factor as $\ln B=\ln\mathcal{Z}_{\rm model}-\ln\mathcal{Z}_{\rm featureless}$. In our analysis, a negative $\ln B$ means that the model is statistically disfavored compared to the featureless case, and a positive value means that the model is favored over the featureless case.

### v2

@@ 1 Introduction
Primordial features are examples of departures in the potential or kinetic term from SR that induce strongly scale-dependent components in the primordial density perturbations. Besides revealing the shape of the inflaton potential (e.g. [19, 20, 21, 22, 23]), primordial features can momentarily increase the energy scale of inflation, potentially boosting the energy reach [24] of this “cosmological collider” [25, 26] by orders of magnitude, allowing for exploration of fundamental physics at even higher energies beyond what can be observed in ground-based collider experiments. Primordial features induced by oscillating massive fields could also be used to probe the background evolutionary history of the primordial universe, providing a possible form of direct evidence for the inflation scenario [27, 28, 29, 30]. (See [31, 32, 33, 34] for reviews on primordial features.) While there is no evidence for such features in current CMB observations, there are some notable, although not statistically significant, deviations from the concordance model at several multipoles in the measured angular spectra. In particular, deviations at very large scales in the $TT$ power spectrum have been observed by both WMAP and Planck. These so-called “$-\ell$ anomalies” are manifest in the $TT$ spectrum as a suppression of power at $\ell<10$ and a dip at $\ell\simeq 20$, the significance of the former depending on the Galactic mask [35] and the latter being approximately $2\sigma$ away from the $\Lambda$CDM best fit. These anomalies have been explored as possibly being primordial in origin [36, 37, 38, 39, 40, 41]. While these anomalies may be explained by statistical fluctuations, especially when noting the large uncertainties at $\ell<10$ due to cosmic variance, future experiments may determine whether these deviations are significant, and they are therefore worth exploring.
@@
A slew of next generation CMB missions following Planck are under development or taking data. Some notable examples include ACT [42], SPT [43], LiteBIRD [44], Simons Observatory [45], and CMB-S4 [46]. Since Planck has achieved nearly cosmic-variance limited measurements in the temperature anisotropies up to $\ell\sim 2500$, most of the improvement in precision from these missions will be in polarization. LiteBIRD is planned for launch by the 2030s and will cover satellite observations with a focus on polarization at large scales ($\ell\leq 200$). Simons Observatory and CMB-S4 are both ground-based experiments. Simons Observatory began taking data in mid-2024, probing multipoles up to $\ell=4000$. CMB-S4 observations are also planned for the 2030s, probing multipoles up to $\ell=5000$. One of the primary goals of these experiments is to obtain high-precision $B$-mode measurements in order to detect primordial gravitational waves. Since the amplitude of the $B$-mode anisotropies is small compared to that of $E$-mode and temperature anisotropies, these missions require unprecedented precision of $E$-mode measurements, with the $EE$ and $TE$ power spectra projected to being constrained nearly to the level of cosmic variance. (Planck measurements have already achieved precision near the level of cosmic variance in the $TT$ power spectrum.) Since models designed to explain the $-\ell$ anomalies in the $TT$ spectrum also predict correlated features in the $EE$ and $TE$ spectra, future experiments will provide a more accurate test of these models.
@@
In this paper, we test the ability of future CMB experiments to constrain or detect cosmological models beyond $\Lambda$CDM that modify the CMB power spectra at large scales with a particular emphasis on primordial feature models. Current candidates to explain the $-\ell$ anomalies include such primordial feature models, which modify the shape of the scalar primordial power spectrum from inflation, as well as models that alter the late-time evolution of the universe, affecting the integrated Sachs-Wolfe (ISW) effect. These models also predict deviations in the CMB bispectra, as well as in the matter power spectrum and bispectrum from Large Scale Structure (LSS), which could help constrain the models further. As a representative sample, we consider three popular models with distinct large-scale effects in the CMB power spectrum: (1) a primordial feature introduced as a step in the inflationary potential that can produce either a bump or a dip at large scales, used to explain the $\ell\simeq 20$ dip in $TT$; (2) a second primordial feature model introducing a kink in the potential, which can lead to either large-scale suppression or amplification, used to explain the $\ell<10$ suppression in $TT$; and (3) a model of dark energy parameterized by the equation of state $w$ of the dark energy fluid and the speed of sound $c_{s}^{2}$ of its perturbations, modifying the late-time ISW effect and also producing either large-scale suppression or amplification. For completeness, we also consider a fourth model, characterized by a micron-sized fifth spacetime dimension that predicts only amplification of the scalar primordial power spectrum. We use the fully Bayesian pipeline introduced in [47] and later improved in [48]. We first compare the models against real Planck data, providing constraints on their parameters, identifying best-fit candidates, and assessing their observational evidence compared to the standard cosmological model. We then perform a Bayesian forecast analysis for the $TT$, $EE$, and $TE$ power spectra using the projected sensitivities to temperature and $E$ modes from future experiments. While the main driver of our analysis is large-scale measurements from LiteBIRD, most models aiming to fit the $-\ell$ anomalies produce modifications to the anisotropy spectra at smaller scales as well. Therefore, we complement our forecast with projected smaller-scale measurements. While such high-quality anisotropy measurements will soon be available from the Simons Observatory, our purpose is to set the most optimistic prospects by including small-scale data, so we provide projected forecasts from CMB-S4. As a final test of LiteBIRD’s capabilities, we ignore the short-scale effects of the primordial feature models and only focus on the suppression and dip anomalies at large scales, which we represent with non-physical, diagnostic templates that produce no small-scale modifications to the primordial power spectrum. We repeat our data and forecast analyses on these templates. We aim to answer three major questions with our analysis: (1) If the true model of the universe is featureless (i.e., consistent with the concordance model), can we hope to rule out large-scale features as statistical fluctuations? (2) Conversely, if the true model contains some kind of large-scale feature, can we detect such a feature? (This question has also been studied for the step model in [41].) (3) In the case of (2), can we further discriminate between different types of features with future observations?
@@ 2 The models
We begin by presenting an overview of the models used in this study. We consider three types of large-scale deviations from the concordance model: a dip, suppression, and an amplification in the $TT$ power spectrum. Our goal is to investigate if future experiments are able to detect such deviations and discriminate between them.
@@
This model is convenient because it allows us to easily relate the free parameters of the model to the scalar power spectrum amplitude $A_{s}$ and tilt $n_{s}$ as well as to the tensor-to-scalar ratio $r$, all defined at a given scale with the Hubble crossing condition $k_{0}=a_{0}H_{0}$
@@
From these equations, we see that $\epsilon_{0}$ controls the amplitude of the tensor-to-scalar ratio $r$. Since our analysis in this paper does not make use of $B$-mode data, we follow previous works [47, 50, 51, 48] and fix $\epsilon_{0}=10^{-7}$ ($r\approx 10^{-6}$) in order not to produce observable $B$ modes.
Dip at Large Scales. We first consider a sharp primordial feature introduced as a bump or dip in the primordial power spectrum. Such a feature can be generated by a “step” in the inflaton potential. The step model was originally motivated by a supergravity model in which several stages of spontaneous symmetry breaking during inflation trigger a series of steps in the inflaton potential [20, 21], and could have other origins as well. After WMAP, such a model is used to fit the $\ell\simeq 20$ dip observed in the $TT$ angular power spectrum [36, 37, 38, 52, 39, 53, 40, 41]. Phenomenologically, the inflaton potential of this model may be written in the following form [21],
@@
where $A$ is the height of the step, $\phi_{0}$ is the feature location, and $\Delta$ is the feature width. The density perturbation in the power spectrum at the feature location can be described as either a dip if $A>0$ (a downward step) or a bump if $A<0$ (an upward step). This model introduces three model parameters: $A$, $\phi_{0}$, and $\Delta$. We relate these model parameters to the following three “effective parameters,” which are more intuitive and more closely related to observables: the maximum deviation in the primordial power spectrum from the featureless spectrum, $\Delta\mathcal{P}/\mathcal{P}_{0}$; the number of $e$-folds from the beginning of inflation until the inflaton encounters the sharp feature, $N_{0}$; and the feature duration in terms of number of $e$-folds, $\Delta N$. The relations between the model and effective parameters are derived in Appendix A.
Large-Scale Primordial Suppression. Next, we consider a primordial feature model that results in large-scale suppression by introducing a “kink” in the inflaton potential. The kink model was originally engineered [19] as an attempt to reconcile a non-$\Lambda$CDM (matter-only) model with the observed galaxy correlation functions. We will use this model phenomenologically despite the original motivation. The potential of the kink model may be written in the following form,
@@
where $A$ is the initial slope of the potential, $\phi_{0}$ is the feature location, and $\Delta$ is the width of the kink. This feature is characterized by a sharp change in the shape of the potential, which is of the form $V\sim V_{0}(\phi)[1+A(\phi_{0}-\phi)]$ for $\phi\ll\phi_{0}$ and $V\sim V_{0}$ for $\phi\gg\phi_{0}$. This model can result in either suppression in the power spectrum if $A>0$ (a negative initial slope) or amplification if $A<0$ (a positive initial slope). This model is therefore motivated by fitting the $\ell<10$ suppression in the $TT$ angular power spectrum. The same model parameters introduced in the step model also describe the kink model, and these are similarly related to the effective parameters $\Delta\mathcal{P}/\mathcal{P}_{0}$, $N_{0}$, and $\Delta N$. We note that other models exist that predict a similar suppression of power that could fit the $-\ell$ suppression observed in the temperature data. For example, double inflation models where inflation consists of two stages driven by two kinetically coupled scalar fields produces a feature signal similar to that of the kink model [54]. Wiggly Whipped Inflation (WWI) [40] combines a step-like feature at intermediate scales with large-scale suppression [55] arising from phase transitions during inflation [^1] A forecast with the WWI model using a Cosmic ORigins Explorer (CORE) like survey was done in [56].. Other single field models such as “punctuated inflation” [57, 58], fast roll initial conditions on the inflaton [59, 55, 60, 61], inflationary models with non-zero spatial curvature [62, 63], or pre-inflationary physics [35, 64, 65] also produce suppression, although more drastic than that considered here.
@@
Finally, while our analysis focuses on the CMB and the ability of this model to fit the $-\ell$ anomalies, it is important to note that models with a time-evolving equation of state also significantly impact the background evolution of the universe and the late-time clustering of matter. Consequently, they are strongly constrained by other datasets, such as supernovae and LSS data. In particular, the equation-of-state parameters are degenerate with other background parameters such as the matter, dark energy, and curvature densities. Late-time observations help in constraining the equation of state further by breaking certain degeneracies. However, since we analyze the constraints and forecasts for CMB-only observations, we do not consider additional datasets. Therefore, the constraints on this model presented here should be considered conservative and are expected to become more stringent with the inclusion of these additional datasets.
@@
where $\epsilon$ and $\delta$ are respectively the first and second SR parameters. We have defined the variable $x\equiv\pi kR_{0}$, where $R_{0}$ is the compact dimension radius at the beginning of inflation with the scale factor normalization chosen in [77]. Since the initial value of $R_{0}$ is not fixed in [78], we convert it into a parameter to be determined by our analysis. We relate $R_{0}$ to the effective parameter $N_{\rm DD}-N_{*}$ via
@@
Figure 1: Examples of large-scale effects in the primordial, $TT$, $TE$, and $EE$ power spectra [from left to right] for each of the models considered in this paper. For these examples, the step model parameters are $\ln(10^{10}A_{s})=3.14$, $n_{s}=0.966$, $N_{0}-N_{*}=-4.0$, $\Delta N=0.5$, and $\Delta\mathcal{P}/\mathcal{P}_{0}=-0.6$. The kink model parameters are the same as those for the step model except for $\Delta N=0.1$ and $\Delta\mathcal{P}/\mathcal{P}_{0}=-0.5$. The featureless and $w-c_{s}^{2}$ models were calculated using the same parameters for the kink model but with $\Delta\mathcal{P}/\mathcal{P}_{0}=0$. The additional $w-c_{s}^{2}$ model parameters are $w=-0.5$ and $c_{s}^{2}=0.01$, with $H_{0}=53.0~{}{\rm km}~{}{\rm s}^{-1}~{}{\rm Mpc}^{-1}$. The dark dimension model parameters are $\ln(10^{10}A_{s})=3.04$, $n_{s}=0.966$, and $N_{\rm DD}-N_{*}=-4.0$ with $A_{s}$ and $n_{s}$ calculated at $k_{*}=0.05~{}{\rm Mpc^{-1}}$. These example parameters were chosen to demonstrate the general characteristics of the $-\ell$ anomalies in $TT$, and observable large-scale amplification for the dark dimension model.
@@ 3 Methodology
We first constrain the effective parameters using the latest CMB temperature and $E$-mode polarization data from the Planck mission [1]. For the multipole range $\ell=2-29$, we use the commander_dx12_v3_2_29 (lowT) likelihood for the temperature anisotropies and the simall_100x143_offlike5_EE_Aplanck_B (lowE) likelihood for the $E$-mode anisotropies [^2] The Planck likelihoods are available on the Planck Legacy Archive: https://pla.esac.esa.int/pla/. To compute the likelihoods, we use the Planck Likelihood code clik 16.0, which available at https://github.com/benabed/clik.. For small angular scales $\ell=30-2500$, we use the 2020 CamSpec v12.5333https://people.ast.cam.ac.uk/~stg20/camspec/ (EG20) likelihood [79]. Because of the wide range of possible effective parameters that could be chosen given the CMB data, especially considering the large uncertainties at low $\ell$, we expect multimodal posterior distributions in our samples. Therefore, we use nested sampling to explore the parameter space using CosmoChord, a PolyChord [^4] https://github.com/PolyChord add-on of CosmoMC [^5] https://cosmologist.info/cosmomc/ [80, 81, 82].
@@
Our choice of ranges for the priors of the cosmological and effective parameters are listed in Table 1. We purposefully choose a prior volume large enough to account for a wide range of possible effects, including the possibility of a featureless spectrum so that the baseline model is also nested into the extended ones. As mentioned, we set the energy scale of inflation to $\epsilon_{0}\equiv 10^{-7}$. For the featureless, step, kink, and $w-c_{s}^{2}$ models, we calculate the primordial power spectrum numerically using BINGO [83]. By convention, we set the end of inflation to $N_{\rm end}=68$ and the pivot scale to exit the horizon at $N_{*}=N_{\rm end}-50=18$ $e$-folds from the onset of inflation. We also define $A_{s}$ and $n_{s}$ at the scale $k_{0}$, which exits the horizon $N_{0}$ $e$-folds from the start of inflation, corresponding to the feature scale for the step and kink models. For the featureless and $w-c_{s}^{2}$ models, we set $\Delta\mathcal{P}/\mathcal{P}_{0}=0$ so that $N_{0}-N_{*}$ and $\Delta N$ are arbitrary, as no features show up in $\mathcal{P}(k)$. We fix them to $N_{0}-N_{*}=-4.0$ and $\Delta N=1.0$. For the dark dimension model, the primordial power spectrum is evaluated directly using Eq. (2.9). For this model, $A_{s}$ and $n_{s}$ are defined at the pivot scale $k_{*}\equiv 0.05~{}{\rm Mpc}^{-1}$, which exits the horizon at $N_{*}=N_{\rm end}-50$. The compact dimension scale $k_{\rm DD}$ exits the horizon at $N_{\rm DD}$ $e$-folds from the onset of inflation. [^6] Note that the dark dimension scale $k_{\rm DD}$ is different from what we call the feature scale $k_{0}$ for the other primordial models we consider. Specifically, $k_{\rm DD}$ does not correspond to the point where the transition from $n_{s}\sim 0$ to $n_{s}\sim 1$ takes place, which one could associate with the feature scale $k_{0}$. Instead, we note that $k_{0}\approx 5k_{\rm DD}$, or $N_{0}-N_{*}=\ln 5+N_{\rm DD}-N_{*}$, which arises from the $\coth x+x{\rm csch}^{2}x$ term in Eq. (2.9). We choose our prior range for $N_{\rm DD}-N_{*}$ such that $N_{0}-N_{*}$ in all of our primordial models has the same prior range.
To determine the improvement in the likelihood of one model over another, we report the Bayes factor for each model, defined as the ratio of evidences between two models. We calculate the Bayes factor as $\ln B=\ln\mathcal{Z}_{\rm model}-\ln\mathcal{Z}_{\rm featureless}$. In our analysis, a negative $\ln B$ means that the model is statistically disfavored compared to the featureless case, and a positive value means that the model is favored over the featureless case. We interpret the Bayes factors using the Jeffreys’ scale [84].